### v1

@@ Abstract
Hyperspectral image (HSI) fusion addresses the challenge of reconstructing High-Resolution HSIs (HR-HSIs) from High-Resolution Multispectral images (HR-MSIs) and Low-Resolution HSIs (LR-HSIs), a critical task given the high costs and hardware limitations associated with acquiring high-quality HSIs. While existing methods leverage spatial and spectral relationships, they often suffer from limited receptive fields and insufficient feature utilization, leading to suboptimal performance. Furthermore, the scarcity of high-quality HSI data highlights the importance of efficient data utilization to maximize reconstruction quality. To address these issues, we propose HyFusion, a novel framework designed to enhance the receptive field and enable effective feature map reusing, thereby maximizing data utilization. First, HR-MSI and LR-HSI inputs are concatenated to form a quasi-fused draft, preserving complementary spatial and spectral details. Next, the Enhanced Reception Field Block (ERFB) is introduced, combining shifting-window attention and dense connections to expand the receptive field, effectively capturing long-range dependencies and reusing features to reduce information loss, thereby boosting data efficiency. Finally, the Dual-Coupled Network (DCN) dynamically extracts high-frequency spectral and spatial features from LR-HSI and HR-MSI, ensuring efficient cross-domain fusion. Extensive experiments demonstrate that HyFusion achieves state-of-the-art performance in HR-MSI/LR-HSI fusion, significantly improving reconstruction quality while maintaining a compact model size and computational efficiency. By integrating enhanced receptive fields and feature map reusing, HyFusion provides a practical and effective solution for HSI fusion in resource-constrained scenarios, setting a new benchmark in hyperspectral imaging. Our code will be publicly available.
@@ I Introduction
One widely adopted solution to address these limitations is to capture high-resolution multispectral images (HR-MSIs) alongside low-resolution hyperspectral images (LR-HSIs). Fusion-based super-resolution (SR) techniques [2, 3, 4] combine HR-MSIs and LR-HSIs to reconstruct high-resolution hyperspectral images (HR-HSIs). These techniques have proven effective in overcoming resolution limitations and maximizing the utility of limited HSI data [5, 6, 7, 8, 9].
Despite their effectiveness, HSI fusion tasks remain inherently challenging, as they require the seamless integration of spatial and spectral information while preserving fine spectral details and enhancing spatial resolution. Existing methods often face two major challenges. First, the loss of critical spatial and spectral information as network depth increases. Second, the underutilization of rich spatial-spectral relationships in the input data. This information loss is particularly problematic in HSI applications, where making the most of the limited available data is crucial due to the scarcity of high-quality HSI datasets. While deep learning (DL)-based HSI fusion methods [10, 11, 12, 13] have advanced the field, many state-of-the-art techniques fail to fully address key challenges such as receptive field expansion and feature map reuse. Consequently, these methods often fall short in maximizing data utilization, which compromises their overall fusion performance. Furthermore, the computational and hardware limitations inherent to HSI applications demand solutions that are not only highly effective but also efficient and easy to implement.
To overcome these challenges, we propose HyFusion, a novel framework designed to enhance the receptive field and reuse feature maps, enabling more efficient data utilization and improved HR-HSI reconstruction. The receptive field plays a crucial role by capturing long-range dependencies and spatial-spectral correlations, as illustrated in Figure 1. While feature map reuse, facilitated by dense connections, retains and amplifies extracted information throughout the network. At the core of HyFusion is the Enhanced Reception Field Block (ERFB), which integrates Improved Swin Transformer Layers (ISTL) with dense connections to expand the receptive field and preserve critical spatial-spectral features, addressing the limitations of existing methods by mitigating information loss and enhancing feature retention. Additionally, HyFusion introduces a Dual-Coupled Network (DCN) comprising two specialized networks: SpeNet, for extracting spectral features, and SpaNet, for extracting spatial features. To guide network convergence, we employ task-specific spectral-aware and frequency-aware loss functions, which align reconstructed HR-HSIs with ground-truth in both frequency and spatial-spectral domains, ensuring enhanced fidelity and minimizing artifacts in the final reconstructions.
@@
  • We propose the HyFusion framework, which emphasizes Enhanced Reception Field and Feature Map Reusing through the proposed ERFB, enabling the network to maximize data utilization by capturing long-range dependencies and reusing extracted features across layers.
  • The proposed DCN dynamically integrates spectral and spatial features from HR-MSI and LR-HSI, leveraging dense connections to fully exploit limited data while maintaining a simple and efficient network architecture.
@@
By integrating enhanced receptive fields and feature map reusing into the fusion process, HyFusion effectively addresses the challenges of data scarcity and information loss, establishing a new benchmark for efficient and high-quality HSI fusion in resource-constrained environments.
@@ II Proposed Method
II-A Architecture of Proposed HyFusion
To design an efficient LR-HSI/HR-MSI fusion model, we aim to use dense-connections [16] to minimize the loss of valuable spectral and spatial information during fusion and enhance the receptive field, thereby ensuring effective information complementarity and high-quality HR-HSI reconstruction. However, cross-domain information fusion requires effective interaction while considering the high computational and hardware resource consumption due to the large size of HSI. Thus, the network architecture must be simple, efficient, and effective for practical applications.
Inspired by ensemble learning in traditional machine learning, which enhances classification performance by using two simple independent classifiers for different modalities and combining their results through voting or weighted averaging, our DCN leverages these advantages. Specifically, we fully exploit the input’s spectral and spatial information by designing two distinct subnetworks to aggregate the information, thereby improving fusion performance with lower computational and spatial complexity.
@@ II-B Proposed Dual-Coupled Network
After obtaining these preliminary drafts, shallow feature extraction is performed using $3\times 3$ convolution with LeakyReLU activation (negative slope $0.2$). The ERFB is then used for deep feature extraction, with skip-connections enhancing feature quality and stabilizing training. The reconstructed HR-HSI is:
@@ II-C Enhanced Reception Field Block
Our design philosophy centers on a dual-enhancement strategy that synergistically combines dense connections for thorough feature reuse with an ISTL for enhanced receptive field and long -range dependency modeling. The dense connection structure ensures that each layer can access and build upon all previous feature representations, effectively preventing information loss and vanishing gradients that commonly occur in deep networks. This is particularly crucial for hyperspectral image fusion, where preserving fine spectral details while maximizing the utility of limited data is essential.
@@
$\mathbf{Z}_{j}=\text{ISTL}(\text{Conv}_{3\times 3}(\text{Cat}(\mathbf{Z}_{0},...,\mathbf{Z}_{j-1}))),j=1,2,3,4\vspace{-0.1cm}$ (7)
@@
$\mathbf{Z}_{5}=\text{Conv}_{3\times 3}(\text{ISTL}(\mathbf{Z}_{4}))\vspace{-0.1cm}$ (8)
@@ IV Conclusion
In this paper, we have proposed HyFusion, a novel framework that effectively addresses the challenges of hyperspectral image fusion through two key innovations: the Enhanced Reception Field Block (ERFB) that integrates dense connections with an Improved Swin Transformer Layer (ISTL) for comprehensive feature utilization and stable information propagation, and a Dual-Coupled Network (DCN) architecture that enables efficient cross-domain feature extraction and data utilization for boosting data-scared hyperspectral imagery application. Our experimental results demonstrate that HyFusion achieves state-of-the-art performance in HR-MSI/LR-HSI fusion while exhibiting remarkable data efficiency with limited training samples. The framework’s success in balancing reconstruction quality with computational efficiency makes it particularly valuable for practical hyperspectral imaging applications in resource-constrained environments.

### v3

@@ Abstract
Hyperspectral image (HSI) fusion addresses the challenge of reconstructing High-Resolution HSIs (HR-HSIs) from High-Resolution Multispectral images (HR-MSIs) and Low-Resolution HSIs (LR-HSIs), a critical task given the high costs and hardware limitations associated with acquiring high-quality HSIs. While existing methods leverage spatial and spectral relationships, they often suffer from limited receptive fields and insufficient feature utilization, leading to suboptimal performance. Furthermore, the scarcity of high-quality HSI data highlights the importance of efficient data utilization to maximize reconstruction quality. To address these issues, we propose HyFusion, a novel Dual-Coupled Network (DCN) framework designed to enhance cross-domain feature extraction and enable effective feature map reusing. The framework first processes HR-MSI and LR-HSI inputs through specialized subnetworks that mutually enhance each other during feature extraction, preserving complementary spatial and spectral details. At its core, HyFusion utilizes an Enhanced Reception Field Block (ERFB), which combines shifting-window attention and dense connections to expand the receptive field, effectively capturing long-range dependencies while minimizing information loss. Extensive experiments demonstrate that HyFusion achieves state-of-the-art performance in HR-MSI/LR-HSI fusion, significantly improving reconstruction quality while maintaining a compact model size and computational efficiency. By integrating enhanced receptive fields and feature map reusing into a coupled network architecture, HyFusion provides a practical and effective solution for HSI fusion in resource-constrained scenarios, setting a new benchmark in hyperspectral imaging. Our code will be publicly available.
@@ I Introduction
One widely adopted solution to address these limitations is to capture high-resolution multispectral images (HR-MSIs) alongside low-resolution hyperspectral images (LR-HSIs). Fusion-based super-resolution (SR) techniques [2, 3, 4] combine HR-MSIs and LR-HSIs to reconstruct high-resolution hyperspectral images (HR-HSIs) [5, 6, 7, 8, 9, 10, 11]. These techniques have proven effective in overcoming resolution limitations for various HSI-related applications [12, 13, 14, 15, 16].
Despite their effectiveness, HSI fusion tasks present significant challenges that require sophisticated solutions. The primary challenge lies in seamlessly integrating spatial and spectral information while simultaneously preserving fine spectral details and enhancing spatial resolution. Current methods encounter two critical limitations that impede their performance. The first is the progressive loss of essential spatial and spectral information as network depth increases, compromising the quality of the final reconstruction. The second is the insufficient utilization of the rich spatial-spectral relationships inherent in the input data, leaving valuable information unexploited. These limitations become particularly acute in the context of HSI applications, where the scarcity of high-quality training data makes efficient information utilization paramount. While learning-based HSI fusion methods [5, 6, 7, 8, 9, 10, 11] have significantly advanced the field, even state-of-the-art approaches have yet to fully address fundamental challenges in information loss and data scarcity for HSI-related applications.
To overcome these challenges, we propose HyFusion, a novel dual-coupled framework designed to maximize data utilization and improve HR-HSI reconstruction quality. Our framework employs a Dual-Coupled Network (DCN) architecture that comprises two specialized networks working in tandem: SpeNet for spectral feature extraction and SpaNet for spatial feature extraction. These networks mutually enhance each other during the feature extraction process, enabling effective cross-domain information exchange. At the core of HyFusion is the Enhanced Reception Field Block (ERFB), which integrates Improved Swin Transformer Layers (ISTL) with dense connections. This integration serves two crucial purposes: expanding the receptive field to capture long-range dependencies and preserving critical spatial-spectral features through comprehensive feature map reuse. This design effectively addresses the limitations of existing methods by mitigating information loss and enhancing feature retention. To ensure optimal reconstruction quality, we employ task-specific spectral-aware and frequency-aware loss functions that align reconstructed HR-HSIs with ground-truth data in both frequency and spatial-spectral domains, thereby enhancing fidelity and minimizing artifacts in the final reconstructions.
@@
  • We propose HyFusion, a novel framework designed to enhance receptive fields and maximize feature map reuse through the ERFB, which combines shifting-window attention and dense connections to effectively capture long-range dependencies while reducing information loss, thereby boosting data efficiency.
  • The proposed DCN dynamically extracts high-frequency spectral and spatial features from LR-HSI and HR-MSI, ensuring efficient cross-domain fusion. Each specialized network leverages complementary information from the other input modality to enhance feature extraction quality.
@@ II Proposed Method
II-A Architecture of HyFusion
The receptive field plays a crucial role by capturing long-range dependencies and spatial-spectral correlations, as illustrated in Figure 1. Given the inherent data scarcity in hyperspectral imaging due to hardware limitations and acquisition costs, maximizing the utility of limited available data becomes crucial. To address this challenge, we propose an efficient LR-HSI/HR-MSI fusion model that leverages dense-connections [19] to enable thorough feature reuse throughout the network. This design principle minimizes the loss of valuable spectral and spatial information during fusion while enhancing the reception field, thereby ensuring effective information complementarity and high-quality HR-HSI reconstruction. The feature reuse mechanism is particularly valuable in HSI applications where training data is limited, as it enables the network to make more effective use of each available sample.
Cross-domain information fusion requires effective integration while managing computational and hardware resource demands of large HSI datasets. This necessitates a network architecture that is both simple and efficient. Therefore, Our proposed DCN, which processes LR-HSI and HR-MSI through specialized subnetworks that mutually enhance each other during feature extraction. This coupled design enables the spatial network to leverage spectral information for better pattern recognition, while the spectral network utilizes spatial context to improve feature discrimination. By facilitating this complementary information exchange, DCN enhances representational capacity while maintaining architectural simplicity.
II-B Dual-Coupled Network
@@
After obtaining these preliminary drafts, shallow feature extraction is performed using $3\times 3$ convolution with LeakyReLU activation (negative slope $0.2$). The ERFB is then used for deep feature extraction. The reconstructed HR-HSI is:
@@ II-C Enhanced Reception Field Block
Our design philosophy focus on a dual-enhancement strategy that synergistically combines dense connections for thorough feature reuse with an ISTL for enhanced receptive field and long-range dependency modeling. The dense connection structure ensures that each layer can access and build upon all previous feature representations, effectively preventing information loss as the network depth increases. This is particularly crucial for hyperspectral image fusion, where preserving fine spectral details while maximizing the utility of limited data is essential.
@@
$\mathbf{Z}_{j}=\text{Cat}(\text{Conv}_{3\times 3}(\text{ISTL}(\mathbf{Z}_{0},...,\mathbf{Z}_{j-1}))),j=1,2,3,4\vspace{-0.1cm}$ (7)
@@
$\mathbf{Z}_{5}=0.2\times\text{Conv}_{3\times 3}(\text{ISTL}(\mathbf{Z}_{4}))+\mathbf{Z}_{0}\vspace{-0.1cm}$ (8)
@@ IV Conclusion
In this paper, we have proposed HyFusion, a novel framework that effectively addresses fundamental challenges in hyperspectral image fusion through its dual-coupled network architecture. By designing specialized subnetworks that mutually enhance each other during feature extraction, HyFusion enables efficient cross-domain information exchange between HR-MSI and LR-HSI inputs. The Enhanced Reception Field Block (ERFB) serves as a crucial component, integrating dense connections with an Improved Swin Transformer Layer (ISTL) to maximize feature utilization and expand the receptive field. Our experiments demonstrate that HyFusion achieves state-of-the-art performance across multiple evaluation metrics while maintaining competitive computational efficiency. Particularly noteworthy is the framework’s strong performance with limited training data, showcasing its effectiveness in addressing the inherent data scarcity challenges in hyperspectral imaging.